OpenStudy (kj4uts):

A breeding group of beavers is introduced into a protected area. After t years the number of beavers in the area is modeled by the function N(t)= 54/0.35+0.68^t 1. How many beavers were initially introduced? 2. Estimate the number of beavers after 5 years. 3. Determine the change in the beaver population between t = 5 and t = 10. (Note: All answers are whole numbers.)

OpenStudy (kj4uts):

I think that the answers are (I just want to make sure): 1. 40 2. 109 3. 36.5 or 37 I think would be a whole number

OpenStudy (kj4uts):

@peachpi

OpenStudy (kj4uts):

N(t)= 54/0.35+0.68^0 = 40 N(t)= 54/0.35+0.68^5 = 109

OpenStudy (kj4uts):

I did the question and got partial points and would like to know where I went wrong?

OpenStudy (tkhunny):

You have written \(N(t) = \dfrac{54}{0.35} + 0.68^{t}\). This appears not to be what you intend. Please write clearly.

OpenStudy (anonymous):

yes the first two are correct. For the third, it looks like you found N(10) instead of the calculating the rate of change

OpenStudy (anonymous):

rate of change = [N(10) - N(5)] / (10 - 5)

OpenStudy (kj4uts):

@tkhunny sorry its N(t)= 54/(0.35+0.68^t)

OpenStudy (tkhunny):

So much better. Notation matters. Be careful.

OpenStudy (anonymous):

misread that. The don't want the rate of change, just the change. N(10) - N(5) is what you're asked for.

OpenStudy (kj4uts):

how do I use N(10) - N(5) do I plug something in?

OpenStudy (anonymous):

yes, you plug in 5 and 10 for t. This is what you had above. N(t)= 54/(0.35+0.68^5) = 109 The N(t) should be N(5) because you substituted 5 for t. This is correct N(5)= 54/0.35+0.68^5 = 109.

OpenStudy (anonymous):

So plug in 10 for t to find N(10)

OpenStudy (kj4uts):

145.4979

OpenStudy (kj4uts):

do I subtract the two?

OpenStudy (tkhunny):

That's N(10). What else would you do to find a change or difference?

OpenStudy (kj4uts):

I am not sure? I thought it would be 145-109?

OpenStudy (anonymous):

yes subtract

OpenStudy (kj4uts):

if I just do 145-109=36 before I got 36.5 and rounded it to 37 and I got partial points maybe its just 36

OpenStudy (anonymous):

it seems so

OpenStudy (kj4uts):

ok thank you :)

OpenStudy (anonymous):

you're welcome